OpenStudy (anonymous):

which line is parallel to y=1/4x+2 ? The choices are : a. y=1/4x+5 b. y=-4x+5 c. y=-1/4x+5 d. y=4x+5

OpenStudy (ranga):

Parallel lines have the same slope. What is the slope of the line y=1/4x+2?

OpenStudy (anonymous):

it's 2

OpenStudy (ranga):

In y = mx + b, m is the slope and b is the y-intercept. In y = 1/4x + 2, what is the slope?

OpenStudy (anonymous):

1/4

OpenStudy (ranga):

Yes. What other answer choices has 1/4 as the slope?

OpenStudy (anonymous):

a is the only one

OpenStudy (ranga):

That is your answer.

OpenStudy (anonymous):

thank you very much :)

OpenStudy (ranga):

You are welcome.